# Avalanche Dynamics and the Effect of Straining in Dislocation Systems with Quenched Disorder

### v1

@@ I Introduction
Despite the stochastic and seemingly unpredictable nature of plasticity of heterogeneous materials, several attempts have been made to establish a direct connection between the features of the microstructure and the emergent plastic response [15]. One of the most powerful predictor of the plastic behavior is the the so-called local yield stress, that is, the critical stress at which the material yields during local probing. This descriptor was first showed to be a powerful tool of predicting the loci of plastic events in model amorphous solids [16] and then its applicability was demonstrated in crystalline materials modeled by discrete dislocation systems as well [17]. Numerical studies on metallic glasses revealed that the statistics of local yield stresses strongly depend on the preparation protocol [18]. In the case of crystalline materials it was shown that global plasticity is related to the local yield stresses through the weakest-link principle, however, this weakest-link picture is fundamentally different depending on whether long- or short-range interactions dominate the dislocation dynamics [17]. Namely, in systems dominated by long-range elastic dislocation-dislocation interactions plastic events are spatially extended which leads to a weakest-link behavior of moderate consistency. If, however, a significant extent of short-ranged quenched disorder (e.g., point defects) is introduced, the plastic events get localized and a more rigorous traditional weakest-link picture is realized. This is related to the concept of the wild to mild transition which explains the empirical observation that depending on the interaction involved a material my exhibit very intense or very mitigated acoustic emission fluctuations during deformation [8].
@@
Figure 1: The phase diagram of 2D discrete dislocation systems with quenched disorder. $\sigma_{\mathrm{ext}}$ and $N$ are the external stress and the number of dislocations, respectively. $Q$ is the ratio of the number of point defects and dislocations characterizing the extent of quenched disorder in the system. The markers indicate the four scenarios that are investigated in this paper. The relaxed $Q=0$ system (scenario I) is governed by long-range dislocation-dislocation interactions and the case of dislocation jamming is realized. The $Q=10$ systems are dominated by the short-range point defect - dislocation interactions, that is, it lies in the pinning regime. Besides the relaxed $Q=10$ configurations (scenario II) two other types of $Q=10$ systems were studied. The configurations were first strained close to the phase transition to sustained flow (scenario III). Then, the configurations areunloaded to zero external stress (scenario IV). The errorbars indicate the normalized stress at which the configuration enters the state of sustained plastic flow.
@@ II Simulation methods
The configurations are prepared in the following way. The dislocations (and point defects) are positioned randomly according to a 2D uniform distribution. Then, the configuration is let to relax and find a metastable configuration while no external stress is applied. This relaxation may be followed by subsequent loading as described later. In our work we studied the ensembles of 100 configurations of each type (see examples of relaxed configurations in Fig. 2). The states at which the behavior of the configurations are investigated are summarized in Fig. 1. In total four different scenarios are studied (which are referred to as scenario I, II, III and IV later in the paper).
@@
    Prestrained configurations with point defects (${Q=10}$) that are obtained from scenario III by instantaneously unloading the systems (that is, decreasing the external stress to zero) and letting them to relax. These configurations only differ from scenario II in their prestraining history.
The dynamical behavior and local yield stresses of these four scenarios listed above were studied by examining the avalanches triggered by quasi-static loading with a spatially homogeneous external stress. An avalanche is considered to have started when the mean dislocation velocity exceeds a predefined threshold $v_{\mathrm{thr}}=10^{-4}v_{0}$. $v_{0}=\frac{\mu b^{2}\sqrt{\rho}}{2\pi(1-\nu)B}$ is the unit of velocities in our simulations where $\mu$ is the shear modulus, $b$ is the length of the Burgers vector, $\rho$ is the dislocation density, $\nu$ is the Poisson ratio and $B$ the dislocation drag coefficient characterizing dislocation mobility. When the mean velocity drops again below this threshold, the avalanches are considered to have ceased (see Fig. 5). For more details of the loading and avalanche detection protocol during local yield stress measurements see the corresponding section on local yield stresses and Ref. [23].
@@ III.1 Static correlations
where $\Delta\bm{r}$ is the relative position of two dislocations of the same sign and $\rho_{\mathrm{s}}^{(2)}$ is the two-point density of same-signed dislocations. $d>0$ means that the relative position is more frequent than in a completely random configuration and $d<0$ corresponds to a relative position which is less likely. The correlation maps are shown in Fig. 3. It is obvious that horizontally and especially vertically there is a strong correlation of dislocations (that is, a dislocation favors a position below or above another one of the same sign). We note that the correlation of opposite-signed dislocations was previously shown to be shorter-range [26], therefore, we only focus on the same-signed correlations in this work. In order to characterize the range of static correlations, the spatial dependence of the correlation function was computed in the vertical direction within a narrow cone of central angle of $\varphi=2\cdot\mathrm{tan}^{-1}(1/10)\approx 11.4^{\circ}$. The correlation functions in the vertical direction obeys
@@
where $\alpha\approx 1.5$ (as derived in Ref. [27] using a variational approach) for all four distinct scenarios investigated but the static correlation length $\xi_{\mathrm{s}}$ is affected by the extent $Q$ of quenched disorder and the straining history (see Fig. 4). The correlations are the strongest $Q=0$ systems (scenario I) and the data suggests data the $d_{\mathrm{s}}(\Delta r)$ has no cutoff or the cutoff length is large ($\xi_{\mathrm{s}}\gtrapprox L$). However, long-range correlations are disrupted by the quenched disorder leading to a static correlation length of $\xi\approx L/4$ in scenario II. One could expect that these correlations get stronger and the static correlation length diverges during straining as the system approaches the phase-transition to the sustained flow state. The prefactor of the dislocation-dislocation correlation, however, gets smaller as the system gets severely strained (scenario III) and it remains small as the system is unloaded (scenario IV). Meanwhile, the correlation length indeed seems to diverge due to the straining with no apparent exponential cutoff in $d_{\mathrm{s}}$ in scenarios III and IV.
@@
Figure 4: The two-point correlation $d_{\mathrm{s}}$ of same-signed dislocations in the vertical direction (for the definition see Eq. (3)). $\Delta r/L$ is the distance of dislocations normalized with the simulation box size. The correlation decays with an exponent of 1.5. In scenario II the correlation has an exponential cutoff characterized by the static correlation length $\xi_{\mathrm{s}}\approx L/4$. In the other three scenarios the data suggest that either there is no cutoff or it is much larger ($\xi_{\mathrm{s}}\gtrapprox L$).
@@ III.2 Dynamic correlations
where $\Delta r=|\Delta\bm{r}|$ is the distance of two dislocations. Figure 6 shows that in the critical $Q=0$ case (scenario I) the triggered dislocation clusters are very extended already upon the onset of the avalanche with a low value of $\delta_{\mathrm{trg}}\approx 1.0$ (see inset). In the presence of point defects (scenario II), however, the decay of $\langle v_{i}v_{j}\rangle$ is much faster with $\delta_{\mathrm{tgr}}\approx 1.8$. Perhaps contra-intuitively during straining (scenario III) and approaching the phase of sustained flow the triggered cores are not getting more extended, on the contrary, they are even more localized with $\delta_{\mathrm{tgr}}\approx 2.1$. After unloading (scenario IV) the system again has the $\langle v_{i}v_{j}\rangle$ behavior seen in the relaxed $Q=10$ state.
@@
The displacement correlation $\langle\Delta x_{i}\Delta x_{j}\rangle$ is computed based on the displacements during the first avalanche. It also exhibits an asymptotic powerlaw dependence on $\Delta r$ with an exponent $\delta_{\mathrm{ava}}$. That is, the asymptotic behavior reads as
@@
As it clear from Fig. 7 the picture that emerges about the spatial extension of avalanches only partly matches the one seen for the activation cores. Similarly to the case of triggered cores, the avalanches are also quite extended in dislocation systems without quenched disorder (scenario I) characterized by an exponent $\delta_{\mathrm{ava}}\approx 0.5$. A significantly higher value of exponent of $\delta_{\mathrm{ava}}\approx 1.4$ at the relaxed $Q=10$ case (scenario II) indicates that the introduction of point defects localizes not only the activation cores but the emerging dislocation avalanches as well. This exponent remains high (in fact, even grows) in the prestrained configuration (scenario IV) as well. However, for severely strained systems (scenario III) $\delta_{\mathrm{ava}}\approx 0.7$ shows that as the system approaches the transition to sustained plastic flow, it starts to behave similarly to the critical $Q=0$ case (scenario I) in terms of the spatial extension of the avalanches.
@@
In this paragraph the tendencies in the dynamical correlations are summarized. The exponents $\delta$ of scenario I and II are quite consistent. In the former case the exponents are low ($10<\delta<$ 1) indicating that the dynamically affected regions of the system are quite extended in space. In the latter case the exponent are significantly higher (around $0.8-0.9$ higher in each case) corresponding to a much more localized dynamic behavior. The strained (loaded) $Q=10$ systems (scenario III), however, exhibit a more complex dynamic pattern. Activation cores (where the triggering of the avalanche happens) are still localized (surprisingly even more than in the relaxed systems) but the dynamical correlations in whole avalanches is much longer-ranged (such as in scenario I). This is consistent with what one would expect being close to the phase transition to sustained flow since correlations lengths near phase transition tend the diverge (e.g. near the Curie point in the context of magnetization). The prestrained systems (scenario IV) have similar dynamical correlations to its relaxed counterpart without any straining history (scenario II), in fact, our simulations show that their behavior at the onset of the first avalanche is practically indistinguishable based on distance dependence of $\langle v_{i}v_{j}\rangle$. We note that both for the triggering and for the whole avalanches the powerlaw regime quantified by the exponent $\delta$ was prevalent over the normalized characteristic dislocation spacing $d/L=1/\sqrt{\rho}L=1/\sqrt{N}$. Below $d/L$ typically we obtained roughly constant (distance-independent) velocity and displacement correlations. This indicates that the dislocations that contribute the most to these correlations (i.e. the ones at the very core of the avalanche) typically have comparable velocities irrespective of their actual distance.
@@ III.3 Participation number
The probability density functions $P$ of PN$(v)$ and PN$(\Delta x)$ are shown in Fig. 8. The results clearly show that more dislocations are involved in both the triggering and the whole avalanche in the $Q=0$ systems (scenario I) than in the $Q=10$ ones (scenarios II, III and IV) irrespective of the deformation history. It can be observed that different straining history does not yield a striking difference between the average behavior of $Q=10$ systems in terms of the participation number (neither at triggering nor for the whole avalanches). It is somewhat contradictory with the more pronounced differences that were shown in the case of dynamic correlations. In the following section the structure of the avalanches is studied (on the level of individual dislocations) and it is demonstrated that strained configurations (with short-range quenched disorder) can produce avalanches that are spatially quite extended despite having a relatively small number of dislocations involved (which resolves the above mentioned seeming contradiction).
Figure 8: The probablity density function $P$ of the participation number PN based on velocities (PN$(v)$) or displacements (PN$(\Delta x)$). For the definition see Eqs. (9) and (10). (Top): The PN based on the instantaneous velocities at the onset of the first avalanche. (Bottom): The PN based on the accumulated displacements during the first avalanche. The insets show the mean PN values and their uncertainty estimated with the Jackknife method.
@@ III.4 Avalanche structure
As it was mentioned previously, when comparing the data in terms of participation numbers and spatial dynamic correlations, the results might seem somewhat contradictory in the case of strained $Q=10$ systems (scenario III). Namely, while during the avalanches the asymptotic $\langle\Delta x_{i}\Delta x_{j}\rangle$ behavior of these systems tends towards the same exponent as for the critical $Q=0$ configurations (scenario I), in terms of PN the strained systems are still much closer to the relaxed $Q=10$ case (scenario II) which contains much more localized events than scenario I. That is, while the avalanches in strained configurations show longer-range dynamical correlations the number of dislocations involved certainly do no seem to grow proportionally. In order to address this discrepancy, in the following the plastic events are analyzed on the level of individual dislocations.
@@
In order to statistically quantify the differences of events in the four scenarios a simple algorithm is introduced to detect clusters of active dislocations and classify plastic events into the three casts presented in Fig. 9. To this end, the accumulated dislocation activity during the avalanches is considered. First, all dislocations with a total displacement below 1% of the largest individual dislocation displacement are discarded and only the ones with displacements above this threshold are kept. Then all dislocation that are within the distance of $3d$ (where $d=L/\sqrt{N}$ is the average dislocation spacing) are virtually connected. The connected clusters of dislocations are considered the cores of the avalanche. Finally, the number of clusters consisting of more dislocations than the third of the size of the largest core are counted and is denoted by $N_{\mathrm{c}}$ (for examples see Fig. 9). Additionally, PN$(\Delta x)$ is computed (including all $N$ dislocations). The classes are then determined as follows. If $N_{\mathrm{c}}>2$, the event is categorized to have ‘many cores’. If not, that is if it only has 1 or 2 cores, and PN$(\Delta x)>N/100=10.24$ then the event is classified as having a ‘large core’, otherwise, it has a ‘small core’. We note that while the choice of thresholds for the number of cores and for the participation number is arbitrary, the same conclusions can be drawn within a sensibly wide range of thresholds.
@@
The execution of this classification on the four scenarios reveals the following picture that is visualized in Fig. 11. As it was shown, the behavior of relaxed $Q=0$ and $Q=10$ systems (scenario I and II) is fundamentally different. In $Q=10$ configurations practically all plastic events are quite small while in the $Q=0$ samples their is a wide selection of events ranging from the motion of single dipoles ($PN\approx 2$) up to events well above our threshold of $N/100=10.24$ (see also Fig. 8). It is the same, however, in both cases that the core of the avalanche is typically one spatially compact active region. The picture is more complex as the $Q=10$ configuration is strained and approaches the phase transition to the state of sustained flow (scenario III). While the growth in terms of PN is incremental, the activation of many small and spatially disjoint parts of the sample during the avalanche is significantly (c. four times) more frequent than in the relaxed case. Individual dislocation velocity data suggests that remote dislocation clusters are drawn into the avalanche by subsequent chain-triggering (see Fig. 10 for an example). This explains the conundrum how the long-range dynamical correlations appear in the strained systems even though the size of the avalanches (in terms of the number of participating dislocations) does not increase proportionally at all. It can be also observed that upon unloading $Q=10$ systems (scenario IV) return to the same behavior characterizing their relaxed counterparts (scenario II) despite their pre-deformation history.
@@
It was demonstrated in previous works that at small strains pure (point defect free) dislocation systems (scenario I) and the ones also containing a great extent of short-range quenched disorder (in the form of, e.g., point defects or precipitates, scenario II) belong to two distinct classes [29, 21]. This has been observed here as well: in both scenarios I and II the avalanches have a single compact core, however, while in scenario I localized and extended events coexists, in scenario II practically all avalanches are localized. The current results imply that the avalanche behavior of immensely strained systems with quenched disorder (scenario III) belong to a third, different regime. In these systems there is diversity of events affecting very local and very extended regions of the sample, but even in the latter case the avalanches consist of a small number of dislocations which are, however, from clusters scattered along the specimen. The proposed phase diagram supplemented with the new regime is schematically shown in Fig. 12 with caption briefly summarizing the nature of avalanches in each regime. It should be stressed that the figure is strictly schematic and the actual shapes of different regimes can not be obtained from the current data.
Figure 12: The regimes of dynamic behavior during avalanche activity. At low external stress in systems dominated by long-range dislocation-dislocation interactions avalanches occur due to the unjamming of dislocations in form of compact events that come from a very wide size-range. At low external stress in systems dominated by short-range interactions upon depinning small localized avalanches occur. At larger stresses however, a significant minority of delocalized avalanches appear which consist of small clusters of moving dislocation scattered across a large area. At large enough stress (which is dependent on the extent of short-ranged quenched disorder) the system reaches the state of sustained plastic flow.

### v2

@@ I Introduction
Despite the stochastic and seemingly unpredictable nature of plasticity of heterogeneous materials, several attempts have been made to establish a direct connection between the features of the microstructure and the emergent plastic response [15]. One of the most powerful predictor of the plastic behavior is the so-called local yield stress, that is, the critical stress at which the material yields during local probing. This descriptor was first showed to be a powerful tool of predicting the loci of plastic events in model amorphous solids [16] and then its applicability was demonstrated in crystalline materials modeled by discrete dislocation systems in both 2D [17] and 3D [18] as well. Numerical studies on metallic glasses revealed that the statistics of local yield stresses strongly depend on the preparation protocol [19]. In the case of crystalline materials it was shown that global plasticity is related to the local yield stresses through the weakest-link principle, however, this weakest-link picture is fundamentally different depending on whether long- or short-range interactions dominate the dislocation dynamics [17]. Namely, in systems dominated by long-range elastic dislocation-dislocation interactions plastic events are spatially extended which leads to a weakest-link behavior of moderate consistency. If, however, a significant extent of short-ranged quenched disorder (e.g., point defects) is introduced, the plastic events get localized and a more rigorous traditional weakest-link picture is realized. This is related to the concept of the wild to mild transition which explains the empirical observation that depending on the interaction involved a material my exhibit very intense or very mitigated acoustic emission fluctuations during deformation [8].
@@
Figure 1: The phase diagram of 2D discrete dislocation systems with quenched disorder. $\sigma_{\mathrm{ext}}$ and $N$ are the external stress and the number of dislocations, respectively. $Q$ is the ratio of the number of point defects and dislocations characterizing the extent of quenched disorder in the system. The markers indicate the four cases that are investigated in this paper. The relaxed $Q=0$ system (condition I) is governed by long-range dislocation-dislocation interactions and the case of dislocation jamming is realized. The $Q=10$ systems are dominated by the short-range point defect - dislocation interactions, that is, it lies in the pinning regime. Besides the relaxed $Q=10$ configurations (condition II) two other types of $Q=10$ systems were studied. The configurations were first strained close to the phase transition to sustained flow (condition III). Then, the configurations are unloaded to zero external stress (condition IV). The errorbars indicate the normalized stress at which the configuration enters the state of sustained plastic flow.
@@ II Simulation methods
The configurations are prepared in the following way. The dislocations (and point defects) are positioned randomly according to a 2D uniform distribution. Then, the configuration is let to relax and find a metastable configuration while no external stress is applied. This relaxation may be followed by subsequent loading as described later. In our work we studied the ensembles of 100 configurations of each type (see examples of relaxed configurations in Fig. 2). The states at which the behavior of the configurations are investigated are summarized in Fig. 1. In total four different conditions are studied (which are referred to as conditions I, II, III and IV later in the paper).
@@
    Prestrained configurations with point defects (${Q=10}$) that are obtained from condition III by instantaneously unloading the systems (that is, decreasing the external stress to zero) and letting them to relax. These configurations only differ from condition II in their prestraining history.
The external stress value $\sigma_{\mathrm{ext}}$ was chosen based on preliminary investigations which were conducted to find the flow stress $\sigma_{\mathrm{flow}}$ at which the system enters the regime of sustained plastic flow. To this end, simulations of 100-dislocation systems were performed. $Q$ was varied as $Q=0,1,2,...,9,10$ and an ensemble of 10 individual simulations were considered for each value of $Q$. After a relaxation step (same as it is described above in the case of $N=1024$ configurations) the configurations were loaded slowly with constant stress rate and homogeneous external stress. The value of $\sigma_{\mathrm{flow}}$ was determined by tracking the mean dislocation velocity because when entering the sustained flow regime the mean dislocation velocity does not decay to zero any longer but takes non-zero values with periodic fluctuations due to the PBC. Based on the scale-invariant nature of our discrete dislocation model the obtained flow stress is universal in the sense that $\sigma_{\mathrm{flow}}/\sqrt{N}$ is the same (for a given $Q$) for systems of different size unless $N$ is very small[27, 28]. The mean values of the rescaled flow stresses are indicated by the gray markers in Fig. 1 and the error-bars denote the standard deviation corresponding to configuration-to-configuration variation. The value of $\sigma_{\mathrm{ext}}$ in condition III was chosen based on this prior analysis so the system is strained close to the phase transition but it is still below the regime of sustained plastic flow.
The dynamical behavior and local yield stresses of these four conditions listed above were studied by examining the avalanches triggered by quasi-static loading with a spatially homogeneous external stress. An avalanche is considered to have started when the mean dislocation velocity exceeds a predefined threshold $v_{\mathrm{thr}}=10^{-4}v_{0}$. $v_{0}=\frac{\mu b^{2}\sqrt{\rho}}{2\pi(1-\nu)B}$ is the unit of velocities in our simulations where $\mu$ is the shear modulus, $b$ is the length of the Burgers vector, $\rho$ is the dislocation density, $\nu$ is the Poisson ratio and $B$ the dislocation drag coefficient characterizing dislocation mobility. When the mean velocity drops again below this threshold, the avalanches are considered to have ceased (see Fig. 5). For more details of the loading and avalanche detection protocol during local yield stress measurements see the corresponding section on local yield stresses and Ref. [24].
@@ III.1 Static correlations
where $\Delta\bm{r}$ is the relative position of two dislocations of the same sign and $\rho_{\mathrm{s}}^{(2)}$ is the two-point density of same-signed dislocations. $d>0$ means that the relative position is more frequent than in a completely random configuration and $d<0$ corresponds to a relative position which is less likely. The correlation maps are shown in Fig. 3. It is obvious that horizontally and especially vertically there is a strong correlation of dislocations (that is, a dislocation favors a position below or above another one of the same sign). We note that the correlation of opposite-signed dislocations was previously shown to be shorter-range [29, 30], therefore, we only focus on the same-signed correlations in this work. In order to characterize the range of static correlations, the spatial dependence of the correlation function was computed in the vertical direction within a narrow cone of central angle of $\varphi=2\cdot\mathrm{tan}^{-1}(1/10)\approx 11.4^{\circ}$. The correlation functions in the vertical direction obeys
@@
where $\alpha\approx 1.5$ (as derived in Ref. [31] using a variational approach) for all four distinct conditions investigated but the static correlation length $\xi_{\mathrm{s}}$ is affected by the extent $Q$ of quenched disorder and the straining history (see Fig. 4). The correlations are the strongest in $Q=0$ systems (condition I) and the data suggests that $d_{\mathrm{s}}(\Delta r)$ has a large cutoff length ($\xi_{\mathrm{s}}\gtrapprox L$). However, long-range correlations are disrupted by the quenched disorder leading to a static correlation length of $\xi\approx L/4$ in condition II. One could expect that these correlations get stronger and the static correlation length diverges during straining as the system approaches the phase-transition to the sustained flow state. The prefactor of the dislocation-dislocation correlation, however, gets smaller as the system gets severely strained (condition III) and it remains small as the system is unloaded (condition IV). In the case of conditions III and IV the exponential cutoff cannot be exactly determined due to the scatter, however, the cutoff in $d_{\mathrm{s}}$ is clearly larger than in condition II consistently with the expected divergence at the critical point.
@@
Figure 4: The two-point correlation $d_{\mathrm{s}}$ of same-signed dislocations in the vertical direction [for the definition see Eq. (3)]. $\Delta r/L$ is the distance of dislocations normalized with the simulation box size. The correlation decays with an exponent of 1.5. In condition II the correlation has an exponential cutoff characterized by the static correlation length $\xi_{\mathrm{s}}\approx L/4$. In the other three conditions the data suggest that the cutoff is much larger ($\xi_{\mathrm{s}}\gtrapprox L$) even though it cannot be determined precisely for conditions III and IV. The solid lines indicate fitted functions [see Eq. (4)].
@@ III.2 Dynamic correlations
where $\Delta r=|\Delta\bm{r}|$ is the distance of two dislocations. Figure 6 shows that in the critical $Q=0$ case (condition I) the triggered dislocation clusters are very extended already upon the onset of the avalanche with a low value of $\delta_{\mathrm{trg}}\approx 1.0$ (see inset). In the presence of point defects (condition II), however, the decay of $\langle v_{i}v_{j}\rangle$ is much faster with $\delta_{\mathrm{tgr}}\approx 1.8$. Perhaps contra-intuitively during straining (condition III) and approaching the phase of sustained flow the triggered cores are not getting more extended, on the contrary, they are even more localized with $\delta_{\mathrm{tgr}}\approx 2.1$. After unloading (condition IV) the system again has the $\langle v_{i}v_{j}\rangle$ behavior seen in the relaxed $Q=10$ state.
@@
The displacement correlation $\langle\Delta x_{i}\Delta x_{j}\rangle$ is computed based on the displacements during the first avalanche. It also exhibits an asymptotic power-law dependence on $\Delta r$ with an exponent $\delta_{\mathrm{ava}}$. That is, the asymptotic behavior reads as
@@
As it clear from Fig. 7 the picture that emerges about the spatial extension of avalanches only partly matches the one seen for the activation cores. Similarly to the case of triggered cores, the avalanches are also quite extended in dislocation systems without quenched disorder (condition I) characterized by an exponent $\delta_{\mathrm{ava}}\approx 0.5$. A significantly higher value of exponent of $\delta_{\mathrm{ava}}\approx 1.4$ at the relaxed $Q=10$ case (condition II) indicates that the introduction of point defects localizes not only the activation cores but the emerging dislocation avalanches as well. This exponent remains high (in fact, even grows) in the prestrained configuration (condition IV) as well. However, for severely strained systems (condition III) $\delta_{\mathrm{ava}}\approx 0.7$ shows that as the system approaches the transition to sustained plastic flow, it starts to behave similarly to the critical $Q=0$ case (condition I) in terms of the spatial extension of the avalanches.
@@
In this paragraph the tendencies in the dynamical correlations are summarized. The exponents $\delta$ of conditions I and II are quite consistent. In the former case the exponents are low ($10<\delta<$ 1) indicating that the dynamically affected regions of the system are quite extended in space. In the latter case the exponent are significantly higher (around $0.8-0.9$ higher in each case) corresponding to a much more localized dynamic behavior. The strained (loaded) $Q=10$ systems (condition III), however, exhibit a more complex dynamic pattern. Activation cores (where the triggering of the avalanche happens) are still localized (surprisingly even more than in the relaxed systems) but the dynamical correlations in whole avalanches is much longer-ranged (such as in condition I). This is consistent with what one would expect being close to the phase transition to sustained flow since correlations lengths near phase transition tend the diverge (e.g. near the Curie point in the context of magnetization). The prestrained systems (condition IV) have similar dynamical correlations to its relaxed counterpart without any straining history (condition II), in fact, our simulations show that their behavior at the onset of the first avalanche is practically indistinguishable based on distance dependence of $\langle v_{i}v_{j}\rangle$. We note that both for the triggering and for the whole avalanches the power-law regime quantified by the exponent $\delta$ was prevalent over the normalized characteristic dislocation spacing $d/L=1/\sqrt{\rho}L=1/\sqrt{N}$. Below $d/L$ typically we obtained roughly constant (distance-independent) velocity and displacement correlations. This indicates that the dislocations that contribute the most to these correlations (i.e. the ones at the very core of the avalanche) typically have comparable velocities irrespective of their actual distance. We also note that our supplementary investigation (not shown here) revealed that changing the (un)loading protocol to gradual (un)loading instead of the instantaneous one presented here do not change the general conclusions of the paper.
@@ III.3 Participation number
The probability density functions $P$ of PN$(v)$ and PN$(\Delta x)$ are shown in Fig. 8. The results clearly show that more dislocations are involved in both the triggering and the whole avalanche in the $Q=0$ systems (condition I) than in the $Q=10$ ones (conditions II, III and IV) irrespective of the deformation history. It can be observed that different straining history does not yield a striking difference between the average behavior of $Q=10$ systems in terms of the participation number (neither at triggering nor for the whole avalanches). It is somewhat contradictory with the more pronounced differences that were shown in the case of dynamic correlations. In the following section the structure of the avalanches is studied (on the level of individual dislocations) and it is demonstrated that strained configurations (with short-range quenched disorder) can produce avalanches that are spatially quite extended despite having a relatively small number of dislocations involved (which resolves the above mentioned seeming contradiction).
Figure 8: The probability density function $P$ of the participation number PN based on velocities (PN$(v)$) or displacements (PN$(\Delta x)$). For the definition see Eqs. (9) and (10). (Top): The PN based on the instantaneous velocities at the onset of the first avalanche. (Bottom): The PN based on the accumulated displacements during the first avalanche. The insets show the mean PN values and their uncertainty estimated with the Jackknife method.
@@ III.4 Avalanche structure
As it was mentioned previously, when comparing the data in terms of participation numbers and spatial dynamic correlations, the results might seem somewhat contradictory in the case of strained $Q=10$ systems (condition III). Namely, while during the avalanches the asymptotic $\langle\Delta x_{i}\Delta x_{j}\rangle$ behavior of these systems tends towards the same exponent as for the critical $Q=0$ configurations (condition I), in terms of PN the strained systems are still much closer to the relaxed $Q=10$ case (condition II) which contains much more localized events than condition I. That is, while the avalanches in strained configurations show longer-range dynamical correlations the number of dislocations involved certainly do no seem to grow proportionally. In order to address this discrepancy, in the following the plastic events are analyzed on the level of individual dislocations.
@@
In order to statistically quantify the differences of events in the four conditions a simple algorithm is introduced to detect clusters of active dislocations and classify plastic events into the three casts presented in Fig. 9. To this end, the accumulated dislocation activity during the avalanches is considered. First, all dislocations with a total displacement below 1% of the largest individual dislocation displacement are discarded and only the ones with displacements above this threshold are kept. Then all dislocation that are within the distance of $3d$ (where $d=L/\sqrt{N}$ is the average dislocation spacing) are virtually connected. The connected clusters of dislocations are considered the cores of the avalanche. Finally, the number of clusters consisting of more dislocations than the third of the size of the largest core are counted and is denoted by $N_{\mathrm{c}}$ (for examples see Fig. 9). Additionally, PN$(\Delta x)$ is computed (including all $N$ dislocations). The classes are then determined as follows. If $N_{\mathrm{c}}>2$, the event is categorized to have ‘many cores’. If not, that is if it only has 1 or 2 cores, and PN$(\Delta x)>N/100=10.24$ then the event is classified as having a ‘large core’, otherwise, it has a ‘small core’. We note that while the choice of thresholds for the number of cores and for the participation number is arbitrary, the same conclusions can be drawn within a sensibly wide range of thresholds.
@@
The execution of this classification on the four conditions reveals the following picture that is visualized in Fig. 11. As it was shown, the behavior of relaxed $Q=0$ and $Q=10$ systems (conditions I and II) is fundamentally different. In $Q=10$ configurations practically all plastic events are quite small while in the $Q=0$ samples their is a wide selection of events ranging from the motion of single dipoles ($PN\approx 2$) up to events well above our threshold of $N/100=10.24$ (see also Fig. 8). It is the same, however, in both cases that the core of the avalanche is typically one spatially compact active region. The picture is more complex as the $Q=10$ configuration is strained and approaches the phase transition to the state of sustained flow (condition III). While the growth in terms of PN is incremental, the activation of many small and spatially disjoint parts of the sample during the avalanche is significantly (c. four times) more frequent than in the relaxed case. Individual dislocation velocity data suggests that remote dislocation clusters are drawn into the avalanche by subsequent chain-triggering (see Fig. 10 for an example). This explains the conundrum how the long-range dynamical correlations appear in the strained systems even though the size of the avalanches (in terms of the number of participating dislocations) does not increase proportionally at all. It can be also observed that upon unloading $Q=10$ systems (condition IV) return to the same behavior characterizing their relaxed counterparts (condition II) despite their pre-deformation history.
@@
It was demonstrated in previous works that at small strains pure (point defect free) dislocation systems (condition I) and the ones also containing a great extent of short-range quenched disorder (in the form of, e.g., point defects or precipitates, condition II) belong to two distinct classes [33, 22]. This has been observed here as well: in both conditions I and II the avalanches have a single compact core, however, while in condition I localized and extended events coexists, in condition II practically all avalanches are localized. The current results imply that the avalanche behavior of immensely strained systems with quenched disorder (condition III) belong to a third, different regime. In these systems there is diversity of events affecting very local and very extended regions of the sample, but even in the latter case the avalanches consist of a small number of dislocations which are, however, from clusters scattered along the specimen. The proposed phase diagram supplemented with the new regime is schematically shown in Fig. 12 with caption briefly summarizing the nature of avalanches in each regime. It should be stressed that the figure is strictly schematic and the actual shapes of different regimes can not be obtained from the current data.
Figure 12: The regimes of dynamic behavior during avalanche activity. At low external stress in systems dominated by long-range dislocation-dislocation interactions (that is, with no or small extent of quenched disorder) avalanches occur due to the unjamming of dislocations in form of compact events that come from a very wide size-range. At low external stress in systems dominated by short-range interactions (that is, in the presence of large extent of quenched disorder) upon depinning small localized avalanches occur. At larger stresses however, a significant minority of delocalized avalanches appear in these systems which avalanches consist of small clusters of moving dislocation scattered across a large area. At large enough stress (which is dependent on the extent of short-ranged quenched disorder) the system reaches the state of sustained plastic flow.